Razor:

Remember my last question? We have to do the same thing with this one. (3^2)^3 +3^0

Razor:

@NarutoandSasuke Uncle?

Razor:

@Elsa213

Razor:

Simplify the expression. Express your answer BOTH with a positive exponent and in standard form .. Is what my question says

Elsa213:

\(\bf\color{blueviolet}{(3^2)^3+3^0}\)

Razor:

Pretty much

Elsa213:

\(\bf{(3^2)}\) 3*3=9 \(\bf{9^3}\) 9*9*9=729 \(\bf{729+3^0}\) \(\small\bf{729+3^0=730}\) Answer: \(\bf\Large\color{aqua}{730}\)

Razor:

oh my god THANK YOU ELSA BLESS YOU A MILLION TIMES OVER!!

Elsa213:

Hehe, ofc. ^.^ <3

Elsa213:

Did you understand, though? o:

Razor:

yesh

Elsa213:

Goodie. cx

Elsa213:

This isn't music, though. ;p You added it in the wrong subject which is probably why you weren't receiving much help. o:

Razor:

Oops XD